OpenStudy (anonymous):

I always have trouble with division. What's 3/2t+2=4? Solve for t.

jimthompson5910 (jim_thompson5910):

The equation is \[\Large \frac{3}{2}t+2=4\] right?

OpenStudy (anonymous):

I need to solve it both ways.

jimthompson5910 (jim_thompson5910):

but that's the equation right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[3/2t + 2 =4\] Like this too.

jimthompson5910 (jim_thompson5910):

that's just a different way to represent it

OpenStudy (anonymous):

That's why I mean't to put both equations.

jimthompson5910 (jim_thompson5910):

First step is to subtract 2 from both sides \[\Large \frac{3}{2}t+2=4\] \[\Large \frac{3}{2}t+2-2=4-2\] \[\Large \frac{3}{2}t=2\] what's next?

OpenStudy (anonymous):

Um, that's where I get stuck -.-

jimthompson5910 (jim_thompson5910):

Then multiply both sides by 2 \[\Large \frac{3}{2}t=2\] \[\Large 2*\frac{3}{2}t=2*2\] \[\Large 3t=4\] Then what?